# WaveGuard: Robust Deepfake Detection and Source Tracing via Dual-Tree Complex Wavelet and Graph Neural Networks

### v1

@@ Abstract
Deepfake technology has great potential in the field of media and entertainment, but it also brings serious risks, including privacy disclosure and identity fraud. To counter these threats, proactive watermarks methods have become a research hotspot by embedding invisible signals to build active protection schemes. However, existing methods are vulnerable to watermark destruction under malicious distortions, which leads to insufficient robustness. Moreover, embedding strong signals may degrade image quality, making it challenging to balance robustness and imperceptibility. To address these issues, this paper proposes a proactive watermarking framework called WaveGuard, which explores frequency domain embedding and graph-based structural consistency optimization. In this framework, the watermark is embedded into the high-frequency sub-bands by dual-tree complex wavelet transform (DT-CWT) to enhance the robustness against distortions and deepfake forgeries. By leveraging joint sub-band correlations and utilizing specific sub-band combinations, the framework performs robust extraction for source tracing and semi-robust extraction for deepfake detection. In addition, the framework adopts dense connectivity strategies to promote feature reuse and improve robustness under complex attacks. To enhance imperceptibility, we propose a Structural Consistency Graph Neural Network (SC-GNN), which constructs graph representations to extract structural features from the original and watermarked images, ensuring consistency between them and reducing perceptual artifacts to improve visual quality. Finally, an attention mechanism named Tanh-based Spatial Embedding Attention Module (TSEAM) is designed to further refine global and local features, thereby improving watermark concealment without compromising robustness. Experimental results show that the proposed method performs exceptionally well in face swap and face replay tasks. The code will be published at https://github.com/vpsg-research/WaveGuard.
@@ I Introduction
The rapid development of artificial intelligence generated content (AIGC) has greatly improved the ability of automation and innovation. However, its wide application also brings risks, especially the deepfake technology, which can generate forged media, threaten privacy, encourage identity theft and spread false information[1]. Deepfake generation techniques mainly includes face swapping and face reenactment. Face swapping[2] replaces the identity features of the target into the source image while preserving the background and semantic information. On the other hand, face reenactment[3, 4] manipulates expressions or poses to map dynamic features onto the target image while maintaining identity consistency. Although deepfake has positive application in some fields, its abuse has caused serious social concern. Thus, it is very important to develop effective preventive measures.
@@
Currently, most deepfake detection methods rely on passive detection, which determines the authenticity of images by analyzing anomalous features present in forged content. Typical methods include spatial domain artifact detection, such as color inconsistency, abnormal boundary mixing and irregular facial muscle movement[5, 6]. Other approaches explore frequency-domain features, identifying forgery clues like spectral distortions and encoding residuals[7]. In addition, some work also introduces biometric consistency (such as blink frequency and heartbeat signal) as auxiliary information to enhance the identification ability of forged content[8]. Although these method are effective in content authenticity verification, it usually comes into play after the generation and dissemination of forged content, and lacks effective traceability and interpretability.
@@
Recently, the concept of proactive approach has emerged as a preventive strategy, and its core idea is to embed invisible signals when the original content is released. This approach is mainly divided into proactive distortions[9, 10, 11, 12, 13] and proactive watermarks[14, 15, 16, 17, 18, 19, 20, 21, 22]. The former interferes with the generation process of a deepfake model by embedding invisible adversarial perturbation in the original image, destroying or counteracting the effect of the deepfake model, and making the generated content appear visually abnormal. While this provides early protection, it lacks traceability and verification capabilities. In contrast, the proactive watermarks method is to embed the watermark information in the original content in advance, and then realize the dual functions of traceability and detectability by detecting the absence or change of the watermark. This method ensures copyright protection and source tracing, especially avoids the unexplained problems in the passive deepfake detection results. However, the existing proactive watermarks methods face challenges in robustness, particularly when dealing with complex deepfake attacks. As deepfake technology advances, the forgery techniques have become more subtle and sophisticated, making it increasingly difficult to extract watermarks effectively. Meanwhile, maintaining invisibility is equally important. Improving robustness often requires stronger embedding, which may introduce perceptual artifacts. Conversely, enhancing invisibility will reduce the robustness against distortion. Thus, balancing robustness and invisibility has become a key challenge for proactive watermarks.
@@
To better understand existing approaches for balancing robustness and invisibility, Fig. 1 presents three different categories of deep learning-based watermarking frameworks. The first two cases adopt spatial domain embedding strategy, and hide watermark information by directly modifying image pixel values. Although the spatial domain method is simple and efficient, it still faces significant challenges when confronted with complex attacks. Because pixel modification inevitably damages the embedded watermark, this method makes the spatial watermark highly sensitive to distortion and deepfake attack. Although Case1 and 2 have shortcomings in robustness, they respectively represent two typical strategies to improve the invisibility of watermark. In Case1, Generative Adversarial Network (GAN)[23] generate high-quality watermarked images through adversarial training, which makes the watermark image visually similar to the original image to enhance invisibility. However, GANs often face problems of unstable training and mode collapse, which result in inconsistent and unstable outputs. In Case2, Variational Autoencoder (VAE)[24] extract latent features from both the original and watermarked images and optimize their similarity through a loss function, aiming to enhance the watermarked image quality while preserving natural image structures. However, due to the constraints of latent space, the reconstructed images generated by VAEs are often blurred and the detailed information is easily lost. Especially when different datasets are pretrained, the effect of feature extraction is not satisfactory.
@@
To address the limitations of existing methods in both robustness and invisibility, we propose a dual-channel watermarking framework called WaveGuard. It integrates frequency-domain embedding with a graph neural network-based structural consistency constraint to support robust source tracing and deepfake detection. Aiming at the problem that the local details are often lost and the directional texture disturbance often leads to the invalidation of the watermark, this framework innovatively uses the translation invariance and directional selectivity of the dual-tree complex wavelet transform (DT-CWT)[25] to embed the watermark into the high-frequency sub-band, thus significantly enhancing its robustness to common distortions and Deepfake forgery. In addition, the traceability and detection function of the watermark are realized by using the correlation between the joint sub-bands. By using different sub-band combinations, ensuring robustness for traceability and semi-robustness for detection. To further enhance performance, the framework adopts dense connection strategies to promote efficient feature reuse and improve the robustness of the watermark. Additionally, in order to solve the problem that traditional watermarking methods introduce obvious artifacts in the visual quality of images, this paper designs a structural consistency graph neural network (SC-GNN), as shown in Case 3 in Fig. 1. Unlike VAE, which extracts the latent features of images, SC-GNN captures the spatial correlation of original images and watermarked images by constructing graphical representations. The graph convolution network (GCN) further extracts and optimizes the structural features, which ensures that the watermark image maintains the integrity of the original image and reduces the perceptual differences. To further enhance imperceptibility, the framework employs a Tanh-based Spatial Embedding Attention Module (TSEAM) to preserve global structures and refine local details, thereby improving the visual quality of watermarked images. Our contributions can be summarized as follows:
@@
Fig. 2: Two-level DT-CWT decomposition.
@@ II-A Proactive Distortions
Proactive defense strategy destroys or counteracts the generation effect of a deepfake model by embedding invisible adversarial perturbation in the original image. This method provides preventive protection before the image is released, ensuring that when the image is maliciously used for deepfake, the generated content will appear visually abnormal. For example, Ruiz et al.[9] proposed a confrontational attack method that destroyed the conditional image translation network through invisible perturbations to prevent deepfake face manipulation. Similarly, CMUA[10] significantly reduces the forgery ability of several deepfake models by generating antagonistic perturbations. In addition, TAFIM[11] uses an attention mechanism to fuse specific perturbations and generate perturbations that cannot be identified by the model, thereby preventing image manipulation. Furthermore, Anti-Forgery[12] interferes with the generation of deepfake content by introducing perceptual perturbations into the Lab color space, creating significant artifacts and maintaining robustness against input transformations. At the same time, DF-RAP[13] provides robust image protection by modeling the social network compression mechanism to defend against deepfake. Although proactive defense can interfere with the generation of deepfake at an early stage, thus providing some protection, it still has some key limitations. For example, powerful generation models may bypass or neutralize these perturbations, thus weakening their effectiveness. In addition, proactive distortions lacks tracking ability and post-verification, and it is difficult to judge whether the image has been tampered with after publication.
@@ II-B Proactive Watermarks
Digital watermarking is a technology to protect copyrights, verify integrity, or source tracing by embedding hidden information in digital media. HiDDeN[26] is the first end-to-end trainable digital robust watermarking framework. It uses the encoder-decoder structure of deep neural networks to hide information and achieves efficient recovery under various image distortions, opening a new research direction in the field of digital watermarking. Aiming at the problem of non-differentiable JPEG compression, MBRS[27] further enhances the robustness of the model by randomly switching real JPEG, simulated JPEG, and noise-free layers in small batches. Similarly, CIN[28] integrates a reversible network and an attention module to enhance robustness while maintaining high invisibility and resistance to distortions. To overcome the limitations of traditional single-stage end-to-end frameworks, TSDL[29] employs a two-stage design to strengthen watermark robustness and improve model stability and convergence speed. Further, ARWGAN[30] enhances both robustness and invisibility through feature fusion and attention mechanisms, aiming to achieve low bit error rates while preserving high visual quality. In addition, Ma et al.[31] constructed a robust watermark model using Swin Transformer and Deformable Convolutional Network (DCN), and proposed a pattern transfer noise layer to improve generalization performance in multi-distortion scenarios. For cross-media distortion (such as print scanning, print camera and screen camera, etc.), researchers have proposed various solutions, including LFM[32], Stegstamp [33], RIHOOP [34], PIMoG [35], GDS-WM [36], and WRAP[37]. These methods target different distortion types by designing optimized strategies to ensure stable watermark extraction and verification after physical printing and re-digitization.
@@
In the emerging field of proactive face forensics, existing methods are limited. Yu et al. [14, 15] proposed an artificial fingerprint embedded within generative models for deepfake detection and source tracing, demonstrating strong scalability and robustness against complex noise. FaceSigns[16] differs from traditional robust watermarking by emphasizing robustness under common distortions while selectively failing under malicious distortions, thus achieving more accurate deepfake detection. Similarly, FaceGuard[17] achieves proactive deepfake detection by embedding semi-robust watermarks. Building on these efforts, Sepmark[18] proposed a separate watermarking framework that effectively addresses both common and malicious distortions by combining robust and semi-robust watermarking using a single encoder and separable decoder. At the same time, Wu et al.[19] proposed a new framework that adjusts robust watermarks into adversarial ones to improve detection and tracing capabilities. It also supports passive detection, making it easier to combine proactive and passive detection. Wang et al. [20] proposed a robust watermarking framework based on identity semantics, achieving deepfake detection and traceability with strong robustness and accuracy. In addition, Lampmark[21] uses an untrained pre-defined pattern watermark to actively defend against deepfake, combining structural sensitivity and robustness to achieve efficient detection. With the growing use of multifunctional watermarking, Dual Defense[22] introduced an invisible, robust watermarking method that combines interference and traceability to prevent effective face-swapping outputs and allow copyright tracking.
@@
Overall, although various methods have explored watermark robustness and deepfake detection, there are still challenges in improving adaptability and imperceptibility. Some methods rely on spatial domain embedding and are easily affected by tampering, while others lack effective visual consistency maintenance mechanism. In contrast, our method integrates frequency domain embedding and structural consistency optimization, and achieves the unity of forgery detection and source tracking while taking into account robustness and invisibility.
@@ III-A1 Dual-tree complex wavelet transform
Dual-Tree Complex Wavelet Transform (DT-CWT) is a wavelet transform with translation invariance and limited redundancy, with a redundancy factor of 2. Unlike traditional DWT, DT-CWT decomposes the image into six directional high-frequency sub-bands, corresponding to angles of $15^{\circ}$, $45^{\circ}$, $75^{\circ}$, $105^{\circ}$, $135^{\circ}$, and $165^{\circ}$. This allows it to accurately capture the directional features and details of the image. As shown in Fig. 2, DT-CWT adopts a circular layout to illustrate the directional and hierarchical structure of two-level decomposition. The first-level decomposition splits the image into six high-frequency sub-bands ($U^{H}_{1,d}$, where $d$ takes values from 1 to 6) and one low-frequency sub-band ($U^{L}$ at the center). The second-level decomposition further splits the first-level low-frequency sub-band, generating six new high-frequency sub-bands ($U^{H}_{2,d}$) and one low-frequency sub-band, with the size of each sub-band halved. Blue arrows represent the first-level high-frequency sub-bands, and red arrows represent the second-level high-frequency sub-bands, with consistent directions across both levels. DT-CWT possesses multi-scale and multi-directional analysis capabilities, which precisely align with the need for accurate characterization of texture boundaries and directional details in watermarking. In the face of compression, blurring, and sophisticated deepfake forgeries, its approximate translation invariance offers a stable and reliable embedding space in the frequency domain, thereby enhancing the robustness of watermark signals against fine-grained distortions.

### v3

@@ Abstract
Deepfake technology has great potential in the field of media and entertainment, but it also brings serious risks, including privacy disclosure and identity fraud. To counter these threats, proactive forensic methods have become a research hotspot by embedding invisible watermark signals to build active protection schemes. However, existing methods are vulnerable to watermark destruction under malicious distortions, which leads to insufficient robustness. Moreover, embedding strong signals may degrade image quality, making it challenging to balance robustness and imperceptibility. To address these issues, this paper proposes a proactive watermarking framework called WaveGuard, which explores frequency domain embedding and graph-based structural consistency optimization. In this framework, the watermark is embedded into the high-frequency sub-bands by dual-tree complex wavelet transform (DT-CWT) to enhance the robustness against distortions and deepfake forgeries. By leveraging joint sub-band correlations and utilizing specific sub-band combinations, the framework performs robust extraction for source tracing and semi-robust extraction for deepfake detection. In addition, the framework adopts dense connectivity strategies to promote feature reuse and improve robustness under complex attacks. To enhance imperceptibility, we propose a Structural Consistency Graph Neural Network (SC-GNN), which constructs graph representations to extract structural features from the original and watermarked images, ensuring consistency between them and reducing perceptual artifacts to improve visual quality. Experimental results show that the proposed method performs exceptionally well in face swap and face replay tasks. The code has been published at https://github.com/vpsg-research/WaveGuard.
@@ I Introduction
The rapid development of artificial intelligence generated content (AIGC) has greatly improved the ability of automation and innovation. However, its wide application also brings risks, especially the deepfake technology, which can generate forged media, threaten privacy, encourage identity theft and spread false information[1]. Deepfake generation techniques mainly includes face swap and face replay. Face swap[2] replaces the identity features of the target into the source image while preserving the background and semantic information. Face replay[3, 4] manipulates expressions or poses to map dynamic features onto the target image while maintaining identity consistency. Although deepfake has positive application in some fields, its abuse has caused serious social concern. Thus, it is very important to develop effective detection methods.
@@
Currently, most deepfake detection methods rely on passive detection, which determines the authenticity of images by analyzing anomalous features present in forged content. Typical methods include spatial-domain artifact detection, such as color inconsistency, abnormal boundary mixing and irregular facial muscle movement[5, 6]. Other approaches explore frequency-domain features, identifying forgery clues like spectral distortions and encoding residuals[7]. In addition, some work also introduces biometric consistency (such as blink frequency and heartbeat signal) as auxiliary information to enhance the identification ability of forged content[8]. Although these method are effective in content authenticity verification, it usually comes into play after the generation and dissemination of forged content, and lacks effective traceability and interpretability.
@@
Recently, the concept of proactive approach has emerged as a preventive strategy, and its core idea is to embed invisible signals when the original content is released. This approach is mainly divided into proactive distortions[9, 10, 11, 12, 13] and proactive watermarks[14, 15, 16, 17, 18, 19, 20, 21, 22]. The former interferes with the generation process of a deepfake model by embedding invisible adversarial perturbation in the original image, destroying or counteracting the effect of the deepfake model, and making the generated content appear visually abnormal. While this provides early protection, it lacks traceability and verification capabilities. In addition, proactive distortions may also weaken the usability of deepfake technology in legitimate applications. In contrast, the proactive watermarks method is to embed the watermark information in the original content in advance, and then realize the dual functions of traceability and detectability by detecting the absence or change of the watermark. This method ensures copyright protection and source tracing, especially avoids the unexplained problems in the passive deepfake detection results. However, the existing proactive watermarks methods face challenges in robustness, particularly when dealing with complex deepfake attacks. As deepfake technology advances, the forgery techniques have become more subtle and sophisticated, making it increasingly difficult to extract watermarks effectively. Meanwhile, maintaining invisibility is equally important. Improving robustness often requires stronger embedding, which may introduce perceptual artifacts. Conversely, enhancing invisibility will reduce the robustness against distortion. Thus, balancing robustness and invisibility has become a key challenge for proactive watermarks.
@@
To better understand existing approaches for balancing robustness and invisibility, Fig. 1 presents three different categories of deep learning-based watermarking frameworks. The first two cases adopt spatial domain embedding strategy, and hide watermark information by directly modifying image pixel values. Although the spatial domain method is simple and efficient, it still faces significant challenges when confronted with complex attacks. Because pixel modification inevitably damages the embedded watermark, this method makes the spatial watermark highly sensitive to distortion and deepfake attack. In Case1, Generative Adversarial Network (GAN)[23] generate high-quality watermarked images through adversarial training, which makes the watermark image visually similar to the original image to enhance invisibility. However, GANs often face problems of unstable training and mode collapse, which result in inconsistent and unstable outputs. In Case2, Variational Autoencoder (VAE)[24] extract latent features from both the original and watermarked images and optimize their similarity through a loss function, aiming to enhance the watermarked image quality while preserving natural image structures. However, due to the constraints of latent space, the reconstructed images generated by VAEs are often blurred and the detailed information is easily lost. Especially when different datasets are pretrained, the effect of feature extraction is not satisfactory.
@@
To address the limitations of existing methods in both robustness and invisibility, we propose a dual-channel watermarking framework called WaveGuard. It integrates frequency-domain embedding with a graph neural network-based structural consistency constraint to support robust source tracing and deepfake detection. Aiming at the problem that the local details are often lost and the directional texture disturbance often leads to the invalidation of the watermark, this framework innovatively uses the translation invariance and directional selectivity of the dual-tree complex wavelet transform (DT-CWT)[25] to embed the watermark into the high-frequency sub-band, thus significantly enhancing its robustness to common distortions and Deepfake forgery. In addition, the traceability and detection function of the watermark are realized by using the correlation between the joint sub-bands. By using different sub-band combinations, ensuring robustness for traceability and semi-robustness for detection. To further enhance performance, the framework adopts dense connection strategies to promote efficient feature reuse and improve the robustness of the watermark. Additionally, in order to solve the problem that traditional watermarking methods introduce obvious visual artifacts, this paper designs a structural consistency graph neural network (SC-GNN), as shown in Case 3 in Fig. 1. Unlike VAE, which extracts the latent features of images, SC-GNN captures the spatial correlation of original images and watermarked images by constructing graphical representations. The graph convolution network (GCN) further extracts and optimizes the structural features, which ensures that the watermark image maintains the integrity of the original image and reduces the perceptual differences. Our contributions can be summarized as follows:
@@
The remainder of this paper is organized as follows. Section II reviews the related work on digital watermarking and deepfake detection. Section III presents the proposed WaveGuard framework in detail. Section IV describes the experimental setup and reports the evaluation results. Section V concludes the paper.
@@ II-A Proactive Distortions
Proactive defense strategy destroys or counteracts the generation effect of a deepfake model by embedding invisible adversarial perturbation in the original image. This method provides preventive protection before the image is released, ensuring that when the image is used in deepfake, the generated content will appear visually abnormal. For example, Ruiz et al.[9] proposed a confrontational attack method that destroyed the conditional image translation network through invisible perturbations to prevent deepfake face manipulation. Similarly, CMUA[10] significantly reduces the forgery ability of several deepfake models by generating antagonistic perturbations. In addition, TAFIM[11] uses an attention mechanism to fuse specific perturbations and generate perturbations that cannot be identified by the model, thereby preventing image manipulation. Furthermore, Anti-Forgery[12] interferes with the generation of deepfake content by introducing perceptual perturbations into the Lab color space, creating significant artifacts and maintaining robustness against input transformations. At the same time, DF-RAP[13] provides robust image protection by modeling the social network compression mechanism to defend against deepfake. Although proactive defense can interfere with the generation of deepfake at an early stage, it still has some key limitations. For example, powerful generation models may bypass or neutralize these perturbations, thus weakening their effectiveness. In addition, proactive distortions lacks tracking ability and post-verification, and it is difficult to judge whether the image has been tampered with after publication.
@@ II-B Proactive Watermarks
Digital watermarking is a technology to protect copyrights, verify integrity, or source tracing by embedding hidden information in digital media. HiDDeN[26] is the first end-to-end trainable digital robust watermarking framework. It uses the encoder-decoder structure of deep neural networks to hide information and achieves efficient recovery under various image distortions, opening a new research direction in the field of digital watermarking. Aiming at the problem of non-differentiable JPEG compression, MBRS[27] further enhances the robustness of the model by randomly switching real JPEG, simulated JPEG, and noise-free layers in small batches. Similarly, CIN[28] integrates a reversible network and an attention module to enhance robustness while maintaining high invisibility and resistance to distortions. To overcome the limitations of traditional single-stage end-to-end frameworks, TSDL[29] employs a two-stage design to strengthen watermark robustness and improve model stability and convergence speed. Tang et al.[30]proposed a two-stage reversible watermarking framework that embeds watermarks into pseudo-Zernike moments using adaptive normalization and optimized DC-QIM, achieving strong resistance to distortions while enabling exact image recovery under benign conditions. Similarly, Huang et al.[31] proposed a texture-aware watermarking method that embeds in high-texture regions using an SSIM-guided regression model, achieving robustness to various attacks while maintaining perceptual quality. In addition, Wang et al.[32] proposed a watermarking framework using DCT-SVD embedding and a template-enhanced network to achieve robust synchronization and accurate extraction under geometric distortions. More recently, Zhu et al.[33] proposed a deep watermarking framework combining Lite-HRNet and a dual U-Net encoder, achieving robustness to screen-shooting and printing while preserving imperceptibility via adaptive noise and adversarial training. For cross-media distortion (such as print scanning, print camera and screen camera, etc.), researchers have proposed various solutions, including LFM[34], Stegstamp [35], RIHOOP [36], PIMoG [37], GDS-WM [38], and WRAP[39]. These methods target different distortion types by designing optimized strategies to ensure stable watermark extraction and verification after physical printing and re-digitization.
@@
In the emerging field of proactive face forensics, existing methods are limited. Yu et al. [14, 15] proposed an artificial fingerprint embedded within generative models for deepfake detection and source tracing, demonstrating strong scalability and robustness against complex noise. FaceSigns[16] differs from traditional robust watermarking by emphasizing robustness under common distortions while selectively failing under malicious distortions, thus achieving more accurate deepfake detection. Similarly, FaceGuard[17] achieves proactive deepfake detection by embedding semi-robust watermarks. Building on these efforts, Sepmark[18] proposed a separate watermarking framework that effectively addresses both common and malicious distortions by combining robust and semi-robust watermarking using a single encoder and separable decoder. At the same time, Wu et al.[19] proposed a new framework that adjusts robust watermarks into adversarial ones to improve detection and tracing capabilities. It also supports passive detection, making it easier to combine proactive and passive detection. Wang et al. [20] proposed an identity-perceptual watermarking framework that enables unified deepfake detection and source tracing by embedding content-dependent semantic watermarks. In addition, Lampmark[21] uses an untrained pre-defined pattern watermark to actively defend against deepfake, combining structural sensitivity and robustness to achieve efficient detection. With the growing use of multifunctional watermarking, Dual Defense[22] employs an original-domain feature impersonation strategy to jointly embed adversarial and traceable watermarks, thereby actively disrupting face swap generation and enabling reliable traceability.
@@
Overall, although various methods have explored watermark robustness and deepfake detection, there are still challenges in improving robustness and imperceptibility. Some methods rely on spatial domain embedding and are easily affected by tampering, while others lack effective visual consistency maintenance mechanism. In contrast, our method integrates frequency domain embedding and structural consistency optimization, and achieves the unity of forgery detection and source tracking while taking into account robustness and invisibility.
@@ III-A1 Dual-tree complex wavelet transform
Dual-Tree Complex Wavelet Transform (DT-CWT) is a wavelet transform with translation invariance and limited redundancy and its redundancy factor is 2. Unlike traditional DWT, DT-CWT decomposes the image into six directional high-frequency sub-bands, corresponding to angles of $15^{\circ}$, $45^{\circ}$, $75^{\circ}$, $105^{\circ}$, $135^{\circ}$, and $165^{\circ}$. This allows it to accurately capture the directional features and details of the image. As shown in Fig. 2, DT-CWT adopts a circular layout to illustrate the directional and hierarchical structure of two-level decomposition. The first-level decomposition splits the image into six high-frequency sub-bands ($U^{H}_{1,d}$, where $d$ takes values from 1 to 6) and one low-frequency sub-band ($U^{L}$ at the center). The second-level decomposition further splits the first-level low-frequency sub-band, generating six new high-frequency sub-bands ($U^{H}_{2,d}$) and one low-frequency sub-band, with the size of each sub-band halved. Blue arrows represent the first-level high-frequency sub-bands, and red arrows represent the second-level high-frequency sub-bands, with consistent directions across both levels. DT-CWT possesses multi-scale and multi-directional analysis capabilities, which precisely align with the need for accurate characterization of texture boundaries and directional details in watermarking. In the face of compression, blurring, and sophisticated deepfake forgeries, its approximate translation invariance offers a stable and reliable embedding space in the frequency domain, thereby enhancing the robustness of watermark signals against fine-grained distortions.
Fig. 2: Two-level Dual-Tree Complex Wavelet Transform decomposition.